OpenStudy (anonymous):

Which of the following groups of side lengths would not be a right triangle? a. 0.6, 0.8, 1.0 b. 4√5, 7√5, 8√5 c. 27, 36, 45 d. 5, 5, 5√2 Please help! Thanks x

OpenStudy (joemarket):

B. 4√5, 7√5, 8√5. I used my calculator. If we use the pythag. theorem, A^2+B^2=C^2 8√5 is equal to 320, but (4√5)^2+(7√5)^2= 325

OpenStudy (anonymous):

Thanks so much! That's what I got too:)

OpenStudy (anonymous):

How is D a right triangle? I'm double checking my work, and I'm having trouble proving that one….

OpenStudy (joemarket):

Because 5^2+5^2=50, and the square root of 50 is 5√2, which is equivalent to 7.071067812

OpenStudy (anonymous):

ahhh.. basic math error. Thanks again!

OpenStudy (joemarket):

lol no problem man!